OpenStudy (anonymous):

The height of a cone is two times its base diameter. What is the volume of the cone in terms of its base radius r?

OpenStudy (nirmalnema):

2h=d d=r/2 4h=r volume = 1/3 pi *r^2*h replace h by r/4